OpenStudy (anonymous):

If f(x) = x − 4 and g(x) = 1 over 2x − 2, find f[g(−6)]. 9 3 −9 −3

OpenStudy (mathstudent55):

First find g(-6). Then input that number into f(x)

OpenStudy (anonymous):

is it -9

OpenStudy (mathstudent55):

\( g(-6) = \dfrac{1}{2(-6) - 2} \) \( g(-6) = \dfrac{1}{-14} \) \(g(-6) \ -\dfrac{1}{14} \)

OpenStudy (anonymous):

oh 3

OpenStudy (mathstudent55):

\( f(- \dfrac{1}{14} = -\dfrac{1}{14} - 4 \)

OpenStudy (anonymous):

i got 3.92 from 1/14-4

OpenStudy (mathstudent55):

\( f(- \dfrac{1}{14}) = -\dfrac{1}{14} - 4 \) \( f(- \dfrac{1}{14}) = -\dfrac{1}{14} - \dfrac{56}{14} \) \( f(- \dfrac{1}{14}) = - \dfrac{57}{14} \)

OpenStudy (mathstudent55):

Notice it's -(1/14) - 4, not 1/14 - 4. The answer is a negative number.

OpenStudy (anonymous):

-9 of course

OpenStudy (mathstudent55):

I don't understand it. How did you get -9?

OpenStudy (mathstudent55):

Oh, I get it.